Black Heritage Trail Tour

8/27/92, Thursday 11:00 AM - 2:30 PM Explore the rich 18th Century heritage of African-Americans in Boston. Cancelled in rain

$1.70

Christian Science Center Tour

8/29/92, Saturday 11:00 AM - 2:00 PM Visit the walk-through globe and other attractions at the mother church of Christian Science. Meet at Student Center info desk in rain

free

Mooch Tours

8/29/92, Saturday 5:00 PM - 8:00 PM Dining at its finest!!! Meet people and eats lots of free food. Fraternity dinners are not just for men anymore! Meet at Student Center info desk in rain

free

Museum of Fine Arts Tour

8/30/92, Sunday 11:00 AM - 2:00 PM A must for the artistically inclined. See the world famous Egyptology exibits as well as the works of many artists of renown. Bring your temporary ID to get in free. Meet by Student Center info desk in rain

$1.70 Newbury Street Tour

8/31/92, Monday 11:00 AM - 1:00 PM From the hottest designer boutiques at one end to the world’s largest record store at the other, Boston’s Newbury Street has shops for all tastes. Come see a place like few in the world. Cancelled in rain

free

Downtown Crossing Tour

8/31/92, Monday 3:00 PM - 5:00 PM Come explore Boston’s best shopping area, including the world famous bargin basement and the largest department store in New England. Cancelled in rain

$1.70

Harbor Cruise

9/1/92, Tuesday 1:00 PM - 4:00 PM Grab a camera and come with us on a leisurely cruise around Boston’s picturesque inner harbor. Cancelled in heavy rain

$5.70

Bridge Circuit

9/3/92, Thursday 9:00 AM - 11:00 AM A tour for the fitness conscious. Explore Boston’s most popular jogging trails along the picturesque Charles River. Cancelled in rain

free

The “Where Everybody Knows Your Name” Tour

9/3/92, Thursday 12:00 PM - 3:00 PM You’ve seen it on TV. Now see it for real!! Come to the only place in Boston where “Everybody knows your name.” Cancelled in heavy rain

$1.70 Night Club Tour

9/3/92, Thursday 8:30 PM - 1:00 AM Dance the night away at one of Boston’s hottest night

clubs. You must be over 18. Bring positive ID. Proper dress required.

Meet at Student Center info desk in rain

$8.00

Radio Station Tour

9/4/92, Friday

9:00 AM - 12:30 PM

Tour the “Rock of Boston” and see the inner workings of one of the area’s best radio stations. Meet by Student Center info desk in rain

free

Wellesley College Tour

9/4/92, Friday

12:00 PM - 6:00 PM

Check out the various charms of MIT’s sister college in the suburbs. Tour beautiful Wellesley College. Meet by Student Center info desk in rain

$6.70

Seafood Tour

9/4/92, Friday 6:00 PM - 9:00 PM Come and eat an authentic New England seafood dinner where the fish were caught that morning. Bring money. Cancelled in rain

$1.70 + dinner

Freedom Trail Tour

9/5/92, Saturday 10:00 AM - 1:00 PM See the places where colonials gathered and revolution brewed. Hike the Freedom Trail and relive the days of Samuel Adams and Paul Revere. Cancelled in rain

$1.70 Shop for Your Room Tour

9/5/92, Saturday 2:00 PM - 5:00 PM Come and find everything you need to furnish your new room (sheets, crates, posters, etc.) from some of Cambridge’s best home supply outlets. Cancelled in rain

free

MIT Athena Clusters Tour

9/5/92, Saturday 6:00 PM - 7:00 PM Learn the locations of the least-frequented clusters of Project Athena workstations.

MEETS IN LOBBY 7 free

Harvard Museums Tour

9/6/92, Sunday 11:00 AM - 1:00 PM See the famous glass flowers and other exhibits of these rarely visited jewels. Meet by Student Center info desk in rain

$1.20

John F Kennedy Museum Tour

9/6/92, Sunday 2:00 PM - 6:00 PM Relive the life and times of one of America’s most charismatic and fondly remembered presidents. Meet by Student Center info desk if it rains

$1.70

All tours leave from the front steps of the Student Center